AJ is saving for a trip for his family and as of today has accumulated $450 in a trip fund. AJ's goal is to accumulate $1800, or 4 times this amount to support the trip. Assuming that all of AJ's trip fund is invested to earn an interest rate of 5%, how long, in years and in fractions of a year, will it take AJ to achieve his goal?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1800=450*(1.05)^n

(1800/450)=(1.05)^n

Taking log on both sides;

log 4=n*log (1.05)

n=log 4/log (1.05)

=28.41 years(Approx).